OpenStudy (anonymous):

compute the given functions. a. f+g b. f-g c. fog d. f/g 1. f(x) -x2/x-2; g(x)=2x 2. f(x)=3x/2x-1; g(x)=x+3/2x-1 3. f(x)=sqrt(x); g(x)=x-4

OpenStudy (anonymous):

fog = f(g(x)) f(x)=sqrt(x); g(x)=x-4 f(g(x)) = sqrt g(x) f(g(x)) = sqrt (x-4)

OpenStudy (anonymous):

f+g = sqrt(x)+x-4 f-g=sqrt(x)-x+4

OpenStudy (anonymous):

f/g = sqrtx/(x-4)